OpenStudy (anonymous):

Multiply and simplify 9x^2/15x^5 times 5x^2/45x^3

OpenStudy (abb0t):

\[\frac{ 9x^2 }{ 15x^5 } \times \frac{ 5x^2 }{ 45x^3 }= \frac{3}{5x^3} \times \frac{1}{9x} = \frac{3}{45x^4} = ?\]

OpenStudy (anonymous):

So do i have to simplify 3/45x^4 or is that the answer?

OpenStudy (abb0t):

well, i think i did my math wrong, but reduce everything.

OpenStudy (anonymous):

So are you sure you multiply acrosslike you did?